# Centralized Reward Agent for Knowledge Sharing and Transfer in Multi-Task Reinforcement Learning

### v2

@@ Abstract
Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning by providing immediate feedback through auxiliary informative rewards. Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework that integrates a centralized reward agent (CRA) and multiple distributed policy agents. The CRA functions as a knowledge pool, which aims to distill knowledge from various tasks and distribute it to individual policy agents to improve learning efficiency. Specifically, the shaped rewards serve as a straightforward metric to encode knowledge. This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring meaningful reward signals. We validate the proposed method on both discrete and continuous domains, including the representative meta world benchmark, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
@@ 1 Introduction
Reinforcement learning (RL) has made significant progress across various domains, like robotics (Kober et al.,, 2013), gaming (Lample and Chaplot,, 2017), autonomous vehicles (Aradi,, 2020), signal processing (Luo et al.,, 2024), and large language models (Shinn et al.,, 2023; Ouyang et al.,, 2022). However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration (Ladosz et al.,, 2022). Reward Shaping (RS) has been proven to be an effective technique to address this challenge by providing additional dense and informative rewards (Sorg et al., 2010b, ; Sorg et al., 2010a, ). Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important for its ability to share and transfer knowledge across tasks. In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward entry to distribute knowledge among different tasks. Therefore, we believe that integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task system learning.
@@
Presently, numerous MTRL algorithms involving knowledge transfer have been developed. Policy distillation methods identify and combine the commonalities across different policies (Rusu et al.,, 2016; Teh et al.,, 2017; Parisotto et al.,, 2016; Xu et al.,, 2024; Ma et al., 2024b, ; Ma et al.,, 2023); representation sharing methods extract and disseminate the common features or gradients among agents (Yang et al.,, 2020; D’Eramo et al.,, 2020; Sodhani et al.,, 2021); and parameter sharing methods design architectural modules to reuse parameters or layers across networks (Sun et al.,, 2022; Cheng et al.,, 2023). Despite their potential, these strategies often encounter challenges such as slow adaptation and response to the transferred knowledge, or suffer from delays in reutilizing and comprehending this information. Therefore, leveraging the RS mechanism, which directly adds a metric to the reward function, offers a compelling alternative to address these limitations.
@@
Regarding the RS techniques, not all shaped rewards serve effectively as the medium for knowledge transfer. Specifically, the intrinsic motivation based rewards are typically designed using heuristics to generate task-agnostic signals. Examples include incorporating exploration bonuses (Bellemare et al.,, 2016; Ostrovski et al.,, 2017; Devidze et al.,, 2022), rewarding novel states (Tang et al.,, 2017; Burda et al.,, 2018), and encouraging curiosity-driven behaviors (Pathak et al.,, 2017; Mavor-Parker et al.,, 2022). Although these approaches broaden extensive exploration, they are not directly related to specific tasks and thus lack transferability. Consequently, we focus on another branch of RS methods, the task-contextual rewards, which automatically learn and encode task-specific information, such as the hidden values, contributions of states, or future-oriented insights, that can be shared effectively across various tasks (Ma et al., 2024a, ; Ma et al., 2025b, ; Ma et al., 2025a, ; Mguni et al.,, 2023; Memarian et al.,, 2021).
@@
To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework (Ma et al., 2024a, ), which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL fRAmework (CenRA). The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents. Each policy agent individually learns control behaviors within their respective tasks and shares their experiences with the CRA. The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information. These rewards are then distributed back to the policy agents to augment their original environmental rewards. Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance the knowledge distribution, considering task similarity and agent learning progress, to ensure a system-wide optimal performance. The main contributions of this paper are summarized as follows:
@@
    We propose the CenRA framework targeted to solve MTRL problems. It incorporates a CRA that functions as a centralized knowledge pool, efficiently distilling and distributing valuable information from various tasks to policy agents, and further adapting to new tasks.
@@
    CenRA leverages the advantage of reward shaping techniques to infuse insights via dense rewards. This approach not only provides a direct metric for policy agents to absorb knowledge but also effectively addresses the sparse-reward challenge.
@@
    We introduce an information synchronization mechanism that considers both task similarity and agent convergence progress to balance multi-task learning. This mechanism provides a novel direction for maintaining system equilibrium in MTRL.
@@
    CenRA is validated in both discrete and continuous control MTRL environments, where the extrinsic rewards are sparse. CenRA outperforms several baseline models in terms of learning efficiency, knowledge transferability, and system-wide performance.
@@ 2 Related Work
Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance (Caruana,, 1993). We discuss existing MTRL literature from three main directions:
@@
Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks (Zeng et al.,, 2021). Policy distillation (Rusu et al.,, 2016) is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al., (2017) introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al., (2016) developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan, (2017) incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al., (2019) further proposed an adaptation mechanism to equalize the impact of each task in policy distillation. Additionally, Xu et al., (2020) explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning. Bai et al., (2023) introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks. Mysore et al., (2022) used separate critics for each task to accompany a single actor to integrate their feedback. These methods mitigate gradient interference to an extent, however, balancing the distribution of knowledge across tasks is crucial. Without a careful trade-off, the performance of the entire system could be compromised.
@@
Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem (D’Eramo et al.,, 2020; Devin et al.,, 2017; Hong et al.,, 2021). Sun et al., (2022) used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it. Yang et al., (2020) employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them. He et al., (2024) introduced the Dynamic Depth Routing framework, which dynamically adjusts the use of network modules in response to task difficulty. Sodhani et al., (2021) leveraged task-related metadata to create composable representations. Cheng et al., (2023) and Lan et al., (2023) both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes. Although these methods demonstrate efficacy in learning shared representations, they may struggle to fully capture the complexity of highly diverse tasks. Moreover, adapting shared structures to new tasks typically requires extra design efforts.
@@
Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities. Model-free meta-learning techniques have been proposed to enhance the multi-task generalization (Finn et al.,, 2017). Yang et al., (2017) designed a sharing network structure that allows an agent to learn multiple tasks concurrently. Vuong et al., (2019) introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning. Wan et al., (2020) proposed a transfer learning framework to handle mismatches in state and action spaces. Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks (Chen et al.,, 2018; Yu et al., 2020a, ), while Ammar et al., (2014) developed a consecutive learning policy gradient approach. These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
@@ 3 Preliminaries
RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024a, ) introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL. The operational architecture of ReLara is shown in Figure 1. Within this framework, the original agent is termed as policy agent, while an assistant reward agent is integrated to enrich the feedback mechanism by generating dense, informative rewards. The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states (and potentially the actions of the policy agent) to craft meaningful reward signals. These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
Figure 1: The ReLara framework Ma et al., 2024a .
@@ 4 Methodology
We propose Centralized Reward Agent fRAmework (CenRA) for MTRL, which incorporates a centralized reward agent (CRA) to support simultaneous reinforcement learning agents across multiple tasks. A high-level illustration of the CenRA framework is shown in Figure 2. The CRA is responsible for extracting general task-specific knowledge from various tasks and distributing valuable information to the policy agents by reconstructing their reward models. The detailed methodology for knowledge extraction and sharing is introduced in Section 4.1. Furthermore, to mitigate the potential disparities in the information that each task could contribute, which might lead to an imbalance in knowledge distribution, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2. Finally, the overall framework of CenRA is presented in Section 4.3.
Figure 2: A high-level illustration of the CenRA framework.
4.1 Knowledge Extraction and Share
@@ 4.1.1 Problem Formulation
The centralized reward agent (CRA) is denoted as $\mathcal{A}^{rwd}$ and multiple policy agents are denoted as $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$. Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones. For example, implementing DQN (Mnih et al.,, 2015) for discrete control tasks, while TD3 (Fujimoto et al.,, 2018) or SAC (Haarnoja et al., 2018a, ) for continuous control tasks. Moreover, the policy of CRA $\mathcal{A}^{rwd}$ is $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is $\pi^{pol}_{i}$, respectively.
@@ 4.1.2 Centralized Reward Agent
The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions. Similar to the ReLara framework Ma et al., 2024a , we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks. The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior. Specifically, $\pi^{rwd}$ maps the Cartesian product of the state space and action space, $S\times A$, to a defined reward space, which constrains the rewards to a range of real numbers, $\mathcal{R}=[R_{min},R_{max}]\subset\mathbb{R}$. For simplicity’s sake, we denote the observation of CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$. To distinguish from the environmental reward, the generated reward is termed as knowledge reward, denoted as $r^{knw}$.
We adopt an off-policy actor-critic algorithm to optimize the CRA Konda and Tsitsiklis, (1999). To aggregate and reuse experiences from all policy agents, a concatenated replay buffer $\mathcal{D}=\bigcup_{i=1}^{N}{\mathcal{D}_{i}}$ is constructed, where $\mathcal{D}_{i}$ represents the replay buffer of each policy agent $\mathcal{A}^{pol}_{i}$. Besides, each transition is augmented with the CRA-generated knowledge reward, $r^{knw}$. Specifically, the transition from policy agent $\mathcal{A}^{pol}_{i}$ stored in the replay buffer is defined as $\tau=(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})$. The augmented transition includes all necessary information for optimizing both the CRA and each corresponding policy agent, thus making the concatenated replay buffer a shared resource across the entire framework and minimizing storage overhead.

### v3

@@ Abstract
Reward shaping is effective in addressing the sparse-reward challenge in reinforcement learning (RL) by providing immediate feedback through auxiliary, informative rewards. Based on the reward shaping strategy, we propose a novel multi-task reinforcement learning framework that integrates a centralized reward agent (CRA) and multiple distributed policy agents. The CRA functions as a knowledge pool, aimed at distilling knowledge from various tasks and distributing it to individual policy agents to improve learning efficiency. Specifically, the shaped rewards serve as a straightforward metric for encoding knowledge. This framework not only enhances knowledge sharing across established tasks but also adapts to new tasks by transferring meaningful reward signals. We validate the proposed method on both discrete and continuous domains, including the representative Meta-World benchmark, demonstrating its robustness in multi-task sparse-reward settings and its effective transferability to unseen tasks.
@@ 1 Introduction
Reinforcement learning (RL) has made significant progress across various domains, such as robotics [Kober et al., 2013], gaming [Lample and Chaplot, 2017], autonomous vehicles [Aradi, 2020], signal processing [Luo et al., 2024], and large language models [Shinn et al., 2023, Ouyang et al., 2022]. However, environments with sparse and delayed rewards remain a significant challenge, as the absence of immediate feedback hinders the agent from distinguishing the value of states and leads to aimless exploration [Ladosz et al., 2022]. Reward Shaping (RS) has been proven to be an effective technique for addressing this challenge by providing additional dense and informative rewards [Sorg et al., 2010b, a]. Concurrently, multi-task reinforcement learning (MTRL) is becoming increasingly important due to its ability to transfer knowledge across tasks. In this context, the auxiliary rewards infused with task-specific information in RS offer a straightforward means to distribute knowledge among different tasks. Integrating RS techniques into MTRL is a highly promising and intuitive direction to enhance the efficacy of multi-task learning systems.
@@
Numerous MTRL algorithms for knowledge transfer have been developed. Policy distillation methods identify and combine commonalities across different policies [Rusu et al., 2016, Teh et al., 2017, Parisotto et al., 2016, Xu et al., 2024]; representation sharing methods extract and share the common features or gradients among agents [Yang et al., 2020, D’Eramo et al., 2020, Sodhani et al., 2021]; and parameter sharing methods design architectural modules to reuse parameters or layers across networks [Sun et al., 2022, Cheng et al., 2023]. Despite their potential, these strategies often face slow adaptation to and limited utilization of transferred knowledge. Therefore, leveraging reward shaping, which directly adds a metric to the reward function, offers a compelling alternative to address these limitations.
@@
Regarding reward shaping, not all shaped rewards effectively serve as a medium for knowledge transfer. Specifically, the intrinsic-motivation-based rewards are typically designed using heuristics to generate task-agnostic signals. Examples include incorporating exploration bonuses [Bellemare et al., 2016, Ostrovski et al., 2017, Devidze et al., 2022], rewarding novel states [Tang et al., 2017, Burda et al., 2018], and encouraging curiosity-driven behaviors [Pathak et al., 2017, Mavor-Parker et al., 2022]. Although these approaches encourage broader exploration, they are not directly related to specific tasks and thus lack transferability. Consequently, we focus on another branch of RS methods, task-contextual rewards, which automatically learn and encode task-specific information, such as hidden values, states contributions, or future-oriented insights, that can be effectively shared across various tasks [Ma et al., 2024a, 2025b, 2025a, Mguni et al., 2023, Memarian et al., 2021].
@@
To share task-related knowledge in MTRL via RS techniques, and inspired by the ReLara framework [Ma et al., 2024a], which integrates an assistant reward agent to densify sparse environmental rewards, we propose the Centralized Reward Agent based MTRL fRAmework (CenRA) [^1] The source code is accessible at: https://github.com/mahaozhe/CenRA. The framework consists of two main components: a centralized reward agent (CRA) and multiple distributed policy agents. Each policy agent individually learns control behaviors within its respective tasks and shares its experiences with the CRA. The CRA extracts common knowledge from these experiences and learns to generate dense rewards that are encoded with task-specific information. These rewards are then distributed back to the policy agents to augment their original environmental rewards. Additionally, given that different tasks may contribute variably to the MTRL system, we introduce an information synchronization mechanism to further balance knowledge distribution by considering task similarity and agent learning progress, thereby ensuring system-wide optimal performance. The main contributions of this paper are summarized as follows:
@@
    We propose the CenRA framework to address MTRL problems. It incorporates a CRA that functions as a knowledge pool, efficiently distilling and distributing valuable information from various tasks to policy agents while adapting to new tasks.
@@
    CenRA leverages reward shaping techniques to infuse insights via dense rewards. This approach not only provides a direct signal for policy agents to absorb knowledge but also effectively addresses the sparse-reward challenge.
@@
    We introduce an information synchronization mechanism that considers both task similarity and agent learning progress to balance multi-task learning. This mechanism provides a novel direction for maintaining system equilibrium in MTRL.
@@
    CenRA is validated in both discrete and continuous control MTRL environments with sparse extrinsic rewards. CenRA outperforms baseline models in learning efficiency, knowledge transferability, and system-wide performance.
@@ 2 Related Work
Multi-task reinforcement learning (MTRL) has attracted significant attention recently due to its potential to share knowledge across multiple tasks, thereby improving learning performance [Caruana, 1993]. We discuss existing MTRL literature from three main directions:
@@
Knowledge Transfer methods focus on identifying and transferring task-relevant features across diverse tasks [Zeng et al., 2021]. Policy distillation [Rusu et al., 2016] is a well-studied approach to extract and share task-specific behaviors or representations that many works are built on: Teh et al. [2017] introduced Distral, which distills a centroid policy from multiple task-policies; Parisotto et al. [2016] developed Actor-Mimic, where a single policy is trained to mimic several expert policies from different tasks; while Yin and Pan [2017] incorporated hierarchical prioritized experience replay buffer to select and learn multi-task experiences; Hessel et al. [2019] further proposed an adaptation mechanism to equalize the impact of each task in policy distillation. Additionally, Xu et al. [2020] explored the transfer of offline knowledge to train policies, and further leveraged online learning for fine-tuning. Bai et al. [2023] introduced a dual-phase learning approach, optimizing individual policies while correcting them across multiple tasks. Mysore et al. [2022] used separate critics for each task to accompany a single actor to integrate their feedback. These methods mitigate gradient interference to an extent, however, balancing the distribution of knowledge across tasks is crucial. Without a careful trade-off, the performance of the entire system could be compromised.
@@
Representation Sharing methods explore architectural solutions of reusing network modules or representing commonalities to the MTRL problem [D’Eramo et al., 2020, Devin et al., 2017, Hong et al., 2021, Ma et al., 2024b, 2023]. Sun et al. [2022] used a parameter compositional approach to learn and share a subspace of parameters, allowing policies for various tasks to be interpolated within it. Yang et al. [2020] employed soft modularization to learn foundational policies and utilized a routing network to generate probabilities to combine them. He et al. [2024] introduced the Dynamic Depth Routing framework, which dynamically adjusts the use of network modules in response to task difficulty. Sodhani et al. [2021] leveraged task-related metadata to create composable representations. Cheng et al. [2023] and Lan et al. [2023] both incorporated attention mechanisms: the former employed attention-based mixture of experts to capture task relationships, while the latter used Temporal Attention for contrastive learning purposes. Although these methods demonstrate efficacy in learning shared representations, they may struggle to fully capture the complexity of highly diverse tasks. Moreover, adapting shared structures to new tasks typically requires extra design efforts.
@@
Single-Policy Generalization methods learn a single policy to solve multiple tasks simultaneously or continuously, in the absence of information from prior policies or task-specific details, in which case, the primary goal is to enhance the policy’s generalization capabilities. Model-free meta-learning techniques have been proposed to enhance the multi-task generalization [Finn et al., 2017]. Yang et al. [2017] designed a sharing network structure that allows an agent to learn multiple tasks concurrently. Vuong et al. [2019] introduced a confidence-sharing agent to detect and define shared regions between tasks to support single policy learning. Wan et al. [2020] proposed a transfer learning framework to handle mismatches in state and action spaces. Additionally, several methods focus on overcoming gradient interference to enhance the generalization in various tasks [Chen et al., 2018, Yu et al., 2020a], while Ammar et al. [2014] developed a consecutive learning policy gradient approach. These methods are efficient in saving computational resources, but the generalization ability of the policy may be constrained when faced with out-of-distribution or previously unseen tasks.
@@ 3 Preliminaries
RL with an Assistant Reward Agent (ReLara) [Ma et al., 2024a] introduces a dual-agent framework designed to tackle the challenge of sparse rewards in RL. Within this framework, the original agent is termed as policy agent, while an assistant reward agent is integrated to enrich the feedback mechanism by generating dense, informative rewards. The reward agent, trained as a self-contained RL agent, autonomously extracts hidden value information from the environmental states and the actions of the policy agent to craft meaningful reward signals. These signals significantly improve learning efficiency by providing immediate and pertinent bonuses.
@@ 4 Methodology
We propose the Centralized Reward Agent fRAmework (CenRA) for MTRL, which incorporates a centralized reward agent (CRA) to support multiple reinforcement learning agents across multiple tasks. A high-level illustration of the CenRA framework is shown in Figure 1. The CRA is responsible for extracting general task-specific knowledge from various tasks and distributing valuable information to the policy agents by reconstructing their reward models. The detailed methodology for knowledge extraction and sharing is presented in Section 4.1. Furthermore, to mitigate the potential disparities in the information that each task contributes, which might lead to an imbalance in knowledge distribution, we introduce an information synchronization mechanism by considering two main factors: the similarity of the tasks and the online learning performance of the policy agents, details given in Section 4.2. Finally, the overall framework of CenRA is presented in Section 4.3.
Figure 1: A high-level illustration of the CenRA framework. The centralized reward agent functions as a knowledge repository, distilling information from various tasks and distributing it to individual policy agents to enhance learning efficiency.
4.1 Knowledge Distillation and Distribution
@@ 4.1.1 Problem Formulation
The centralized reward agent (CRA) is denoted as $\mathcal{A}^{rwd}$ and multiple policy agents are denoted as $\{\mathcal{A}^{pol}_{1},\allowbreak\mathcal{A}^{pol}_{2},\dots,\mathcal{A}^{pol}_{N}\}$. Each policy agent $\mathcal{A}^{pol}_{i}$ operates independently to complete its corresponding task $T_{i}$, utilizing appropriate RL algorithms as backbones. For example, implementing DQN [Mnih et al., 2015] for discrete control tasks, while TD3 [Fujimoto et al., 2018] or SAC [Haarnoja et al., 2018a] for continuous control tasks. Moreover, the policy of CRA $\mathcal{A}^{rwd}$ is $\pi^{rwd}$, and the internal policy of policy agent $\mathcal{A}^{pol}_{i}$ is $\pi^{pol}_{i}$.
@@ 4.1.2 Centralized Reward Agent
The CRA $\mathcal{A}^{rwd}$ aims to extract environment-relevant knowledge and distribute it to policy agents by generating additional dense rewards to support their original reward functions. Similar to the ReLara framework [Ma et al., 2024a], we model the CRA as a self-contained RL agent, yet, as an extension to ReLara, our CRA is designed to concurrently interact with multiple policy agents and their respective tasks. The CRA’s policy $\pi^{rwd}$ generates continuous rewards given both an environmental state and a policy agent’s behavior. Specifically, $\pi^{rwd}$ maps the Cartesian product of the state space and action space, $S\times A$, to a defined reward space, which constrains the rewards to a range of real numbers, $\mathcal{R}=[R_{min},R_{max}]\subset\mathbb{R}$. For simplicity, we denote the observation of the CRA as $s^{rwd}=(s_{i},a_{i})$, where $s_{i}\sim T_{i}$ and $a_{i}\sim\pi^{pol}_{i}(s_{i})$. To distinguish from the environmental reward, the generated reward is termed as knowledge reward, denoted as $r^{knw}$.
We adopt an off-policy actor-critic algorithm to optimize the CRA [Konda and Tsitsiklis, 1999]. To aggregate and reuse experiences from all policy agents, a concatenated replay buffer $\mathcal{D}=\bigcup_{i=1}^{N}{\mathcal{D}_{i}}$ is constructed, where $\mathcal{D}_{i}$ represents the replay buffer of each policy agent $\mathcal{A}^{pol}_{i}$. Besides, each transition is augmented with the CRA-generated knowledge reward, $r^{knw}$. Specifically, the transition from policy agent $\mathcal{A}^{pol}_{i}$ stored in the replay buffer is defined as $\tau=(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})$. The augmented transition includes all necessary information for optimizing both the CRA and each corresponding policy agent, thus making the concatenated replay buffer a shared resource across the entire framework and minimizing storage overhead.